OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. 4x2 + ___x + 49

OpenStudy (amistre64):

if its a perfect square trinomial then it is the results of:\[(\sqrt{4x^2}+\sqrt{49})^2\]

OpenStudy (lgbasallote):

for the middle term use \(\small 2\sqrt{(4x^2)(49)}\)

OpenStudy (amistre64):

indeed